# Probing the Bounce Energy Scale in Bouncing Cosmologies with Pulsar Timing Arrays

### v1

@@ A.3 Posteriors
Figures 913 display the one- and two-dimensional posterior distributions for the six conventional SGWB source models: supermassive-black-hole binaries (SMBHBs), inflationary gravitational waves (IGW), cosmic strings (CS), domain walls (DW), first-order phase transitions (FOPT), and scalar-induced GWs (SIGW), following Ref. [45] (c.f. Fig.2 and Fig.7-10 of [45]).
@@ Appendix B Accelerating Bayesian Inference with Normalizing-Flow Machine Learning
In Fig. 14 we show the posterior in the $(\rho_{s\downarrow}^{1/4},w_{1})$ plane obtained from a normalizing-flow–based machine-learning architecture trained on the ten most sensitive pulsars from the NANOGrav 15-yr dataset [99]. Following the methodology of Ref. [130] (see also Refs. [145, 146]), the flow is trained on the analytic SGWB spectrum of the bouncing model (Eq. 1). The normalizing‐flow posterior closely reproduces our MCMC result (Fig. 4, Sec. III) while reducing inference time from $\sim 10$ days to $\sim 20$ hours (see Ref. [130] for technical details), demonstrating a scalable route to rapid Bayesian analyses as PTA datasets—and future arrays such as SKA—expand. Fig. 15 presents a high-resolution view of the “left” branch ($w_{1}\sim 0.3$), highlighting the consistency between MCMC and the normalizing-flow–based approach.
@@
Figure 14: Comparison of posterior distributions in the $(\rho_{s\downarrow}^{1/4},w_{1})$ plane inferred via MCMC (blue dashed contours) and via a normalizing‐flow–based machine‐learning emulator (filled shading), trained on the ten most sensitive pulsars from the NANOGrav 15-yr dataset [99]. Both methods fit Eq. (1). The dashed horizontal line denotes the Planck scale $M_{\rm pl}=G^{-1/2}\simeq 1.22\times 10^{16}\,\mathrm{TeV}$.
@@
Figure 15: High‐resolution view of the “left” posterior branch ($w_{1}\sim 0.3$) from Fig. 14, highlighting the consistency between MCMC (blue dashed) and normalizing‐flow–based machine‐learning (shaded) inferences.

### v4

@@ A.3 Posteriors
Figures 913 display the one- and two-dimensional posterior distributions for the six conventional SGWB source models: supermassive-black-hole binaries (SMBHBs), inflationary gravitational waves (IGW), cosmic strings (CS), domain walls (DW), first-order phase transitions (FOPT), and scalar-induced GWs (SIGW), obtained by employing Ceffyl Lamb:2023jls to perform Bayesian fit of each model (Sec. A.1) to the full PTA dataset ( NANOGrav 15-year, EPTA DR2, PPTA DR3 and IPTA DR2) NANOGrav:2023gor ; EPTA:2023fyk ; Reardon:2023gzh ; Antoniadis:2022pcn ; ZenodoNG ; NANOGrav:2023hde ; ZenodoEPTA ; EPTA:2023sfo ; PPTADR3 ; Zic:2023gta ; IPTADR2 ; Perera:2019sca , following Ref. Li:2024oru (c.f. Fig.2 and Fig.7-10 of Li:2024oru ).
@@ Appendix B Accelerating Bayesian Inference with Normalizing-Flow Machine Learning
In Fig. 14 we show the posterior in the $(\rho_{s\downarrow}^{1/4},w_{1})$ plane obtained from a normalizing-flow–based machine-learning architecture trained on the ten most sensitive pulsars from the NANOGrav 15-yr dataset ZenodoNG . Following the methodology of Ref. Lai:2025xov (see also Refs. Shih:2023jme ; Vallisneri:2024xfk ), the flow is trained on the analytic SGWB spectrum of the bouncing model (Eq. 1). The normalizing-flow posterior closely reproduces our MCMC result (Fig. 4, Sec. 3) while reducing inference time from $\sim 10$ days to $\sim 20$ hours (see Ref. Lai:2025xov for technical details), demonstrating a scalable route to rapid Bayesian analyses as PTA datasets—and future arrays such as SKA—expand. Fig. 15 presents a high-resolution view of the “left” branch ($w_{1}\sim 0.3$), highlighting the consistency between MCMC and the normalizing-flow–based approach.
@@
Figure 14: Comparison of posterior distributions in the $(\rho_{s\downarrow}^{1/4},w_{1})$ plane inferred via MCMC (blue dashed contours) and via a normalizing-flow–based machine-learning emulator (filled shading), trained on the ten most sensitive pulsars from the NANOGrav 15-yr dataset ZenodoNG . Both methods fit Eq. (1). The dashed horizontal line denotes the Planck scale $M_{\rm pl}=G^{-1/2}\simeq 1.22\times 10^{16}\,\mathrm{TeV}$.
@@
Figure 15: High-resolution view of the “left” posterior branch ($w_{1}\sim 0.3$) from Fig. 14, highlighting the consistency between MCMC (blue dashed) and normalizing-flow–based machine-learning (shaded) inferences.
Appendix C A Broken Power-Law Feature for Higher Frequency GW Searches and BBN and CMB Constraints
Besides PTAs, the bouncing-universe SGWB may also be probed by space- and ground-based detectors (e.g., LISA, Taiji, Tianqin, aLIGO + aVirgo + KAGRA, Einstein Telescope, Cosmic Explorer, see Schmitz:2020syl ; Annis:2022xgg ; Bi:2023tib for the summary of their sensitivities, and for recent developments in gravitational wave cosmology incorporating these GW detectors, for example, see Bian:2025ifp ; Yao:2025wfd ; Chen:2023zkb ; Chen:2024mwg and reference therein), while BBN and CMB observations impose bounds on its amplitude over a broad frequency range.
In Fig. 16, we illustrate forecast projections and sensitivities from current and forthcoming GW detectors (colored solid and dashed curves, following Schmitz:2020syl ; Bi:2023tib ); BBN and CMB bounds (red and purple solid horizontal lines, $\Omega_{\mathrm{GW}}^{\mathrm{BBN}}h^{2}(f)\lesssim 7.8\times 10^{-6}$ and $\Omega_{\mathrm{GW}}^{\mathrm{CMB}}h^{2}(f)\lesssim 1.7\times 10^{-6}$, following Smith:2006nka ; Pagano:2015hma ); and four representative examples of the bouncing-universe model (solid pink, orange, green, and blue lines, following Fig. 6.). Without a broken power-law feature, these examples exceed the BBN and CMB limits for $f>10^{-7}\,\mathrm{Hz}$. This implies that the bouncing-universe interpretation likewise requires a (quasi) broken power-law feature (dash-dotted pink, orange, green, and blue lines) at frequencies just above the PTA band to remain consistent with BBN and CMB constraints, analogous to the inflationary interpretation (blue dashed line) and its broken power-law feature (blue dotted line) Kuroyanagi:2020sfw ; Benetti:2021uea . The thin colored dashed curves correspond to six conventional SGWB models: astrophysical SMBHBs (dark red dashed line and gray band Rosado:2015epa ), cosmic strings (green Bian:2023dnv ), domain walls (brown Bian:2022qbh ), first-order phase transitions (blue Bian:2023dnv ), SIGWs (sea-green Bian:2023dnv ; Li:2024lxx ), and non-Gaussianity (SIGW-NG, black Liu:2023ymk ). These models also incorporate a (quasi) broken power-law feature.
Furthermore, Fig. 16 also illustrates that the prospect of probing a bouncing cosmology with space- and ground-based gravitational-wave detectors (such as LISA, Taiji, Tianqin, etc.) depends on the SGWB spectral index above the broken power-law scale. For example, the orange and blue dash-dotted curves—with only a mild red tilt after the break (e.g., $n_{T}\sim 0.25$ for a break at $f\sim 10^{-7}$z)—remain within the sensitivity of these detectors at higher frequencies, whereas the pink and green dash-dotted curves—with a steep red tilt—fall below their sensitivity (e.g., $n_{T}<-2$ for a break at $f\sim 10^{-7}$z).
Figure 16: Forecast projections and sensitivities from current and forthcoming GW detectors (colored solid/dashed curves); BBN and CMB bounds (red/purple solid horizontal lines); and four representative examples of the bouncing-universe model (solid pink, orange, green, and blue lines). Without a broken power-law feature, these examples exceed the BBN and CMB limits for $f>10^{-7}$ Hz. Consequently, the bouncing-universe interpretation likewise requires a (quasi) broken power-law feature (colored dash-dotted lines) at frequencies just above the PTA band to remain consistent with BBN and CMB constraints, analogous to the inflationary interpretation (blue dashed line) and its broken power-law feature (blue dotted line). The thin colored dashed curves correspond to six conventional SGWB models: astrophysical SMBHBs (dark red dashed line and gray band), cosmic strings (green), domain walls (brown), first-order phase transitions (blue), SIGWs (sea-green), and non-Gaussianity (SIGW-NG, black). These models also exhibit (quasi) broken power-law behavior. Furthermore, the prospect of probing a bouncing cosmology with space- and ground-based detectors (e.g., LISA, Taiji, Tianqin, aLIGO+aVirgo+KAGRA, Einstein Telescope, Cosmic Explorer) depends on the SGWB spectral index above the broken power-law scale. For example, the orange and blue dash-dotted cases—with a mild red tilt after the break ($n_{T}\sim 0.25$ for a break at $f\sim 10^{-7}$ Hz)—lie within detector sensitivity, whereas the pink and green dash-dotted cases—with a steep red tilt ($n_{T}<-2$ for a break at $f\sim 10^{-7}$ Hz)—fall below sensitivity at higher frequencies.